If you have 750.0 g of HNO3 and excess Na2CO3, how many grams of carbon dioxide will be produced?


1
Expert's answer
2021-04-13T02:07:32-0400

First we need the balanced equation for the reaction;


2HNO3 + Na2CO3 "\to" 2NaNO3 + H2O + CO3


Moles of HNO3 used "=\dfrac{Mass}{MM} = \dfrac{750.0g}{63.01g/mol} = 11.9029mol"


Since Na2CO3 is in excess, HNO3 is the limiting reagent


Thus moles of CO2 will be determined by moles of HNO3

Mole ratios of HNO3 : CO2 is 2:1


Moles of CO2 "=" "\dfrac{1}{2}x11.9029"


= 5.9514 mol


Mass of CO2 = Moles x MM

= 5.9514 mol x 44.01 g/mol

= 261.9227g